OpenStudy (anonymous):

Find the probability that a non-leap year chosen at random has: a) 52 Sundays b) 53 Sundays

OpenStudy (anonymous):

I say that the answers are: a) P(52 Sundays in a non-leap year ) = 1 (becoz 52 Sundays are guaranteed)

OpenStudy (anonymous):

you are supposed to say 1/7

OpenStudy (anonymous):

b) P(53 Sundays in a non-leap year) = 1/7 (the one extra day can be a Sunday)

OpenStudy (anonymous):

@satellite73 since 52 Sundays in a year are guaranteed so won't the probability be 100% i.e. 1 ??

OpenStudy (anonymous):

by 52sundays they mean exactly 52, as in 52 but not 53

OpenStudy (anonymous):

but this is a totally crap problem because there is no probability in it

OpenStudy (anonymous):

it is like asking what is the probability tomorrow is tuesday. the question is ill posed and makes no sense. tell your teacher that this problem is garbage, a year either contains 52 sundays or 53 and it is not determined by probability

OpenStudy (anonymous):

well if it asked in an exam, one can't say it is a crap problem. one has to answer...

OpenStudy (anonymous):

what it means is, if all the years were put on slips of paper and put in a hat, and you picked out one, what is the probability that that year contained 53 sundays. that question makes sense

OpenStudy (anonymous):

what about the second one b) ??

OpenStudy (anonymous):

anyway your answer is right, 1/7 for 53, 6/7 for 52

OpenStudy (anonymous):

I am not sure about the a) part because every year has 52 Sundays...

OpenStudy (anonymous):

this question has been asked of students before. you can probably google it and you will see that your answer is correct, and probably also find a discussion on why this problem is stupid

OpenStudy (anonymous):

anyway, thanks for your help......

OpenStudy (anonymous):

yes because you are supposed to be a mind reader and know that when they write 52 sundays, they mean exactly 52 sundays, at least that is my guess

OpenStudy (anonymous):

when not so clever math teachers make up probability problems, there is lots of room for ambiguity and error. this is a very good example of that

OpenStudy (anonymous):

well I googled it and some say probability is 1 (sure event) others say it is 6/7

OpenStudy (anonymous):

@satellite73 can't we look at a) as follows P(52 Sundays in a non-leap year) = 1 - p(53 Sundays in a non-leap year) ?? if yes, then P(52 Sundays) = 1 - 1/7 = 6/7